# Resonance RoPE: Improving Context Length Generalization of Large Language Models

### v1

@@ 1 Introduction
${}^{\dagger}$anada CIFAR AI Chair. Corresponding author.
@@
Recent efforts to enhance TSTL performance have focused on LLMs equipped with Rotary Position Embedding (RoPE) (Su et al., 2024), such as LLaMA (Touvron et al., 2023a, b) and Mistral (Jiang et al., 2023), owing to their exceptional capabilities and widespread adoption. These initiatives aim to refine the test-time computation of RoPE position embedding by introducing a scaling factor to either the position index of each token (Chen et al., 2023) or RoPE’s base value (Xiong et al., 2023; Liu et al., 2023b; Peng et al., 2023). These methods ensure that the position embeddings for out-of-distribution (OOD) positions remain within the range experienced during pre-training. This minimizes the need for the model to adapt to new position embedding value ranges, a task that is inherently difficult.
In this paper, we introduce Resonance RoPE, a novel technique designed to further narrow the generalization gap on position embeddings in TSTL scenarios. Recognizing that RoPE’s position embedding is governed by a complex, non-linear function, we posit that minimizing extrapolation on OOD positions, while crucial, is insufficient. We argue that it is equally vital to address the interpolation of RoPE features at the OOD positions. By implementing Resonance RoPE, we effectively eliminate the generalization gap for more than half of the position embedding features in LLaMA and LLaMA2 in TSTL scenarios. Furthermore, our approach is compatible with RoPE and any RoPE-based scaling techniques, enhancing their performance in TSTL situations without the need for additional computational resources during training or inference.
Additionally, to facilitate further research on position embeddings, we present a new synthetic benchmark tailored for TSTL scenarios, named PosGen. Improving position embeddings for TSTL requires a detailed analysis of the cause of failures in handling longer contexts. However, current benchmarks, such as those measuring perplexity in long context (Rae et al., 2020; Huang et al., 2021; Wu et al., 2022) and most synthetic TSTL tasks (Liu et al., 2023a; Kazemnejad et al., 2023) face a common issue: the difficulty of generating the next token increases with context length. This makes it difficult to determine whether a model’s failure is due to its inability to generate more complex tokens or its failure to recognize out-of-distribution (OOD) positions. PosGen addresses this limitation by standardizing the difficulty level of token generation across all positions. This ensures that any observed shortcomings are directly related to the model’s inability to identify and handle new token positions effectively.
@@ 2.1 Scaling of RoPE Position Encoding
Recent efforts in extending LLMs’ context window focus on manipulating position embedding (PE), particularly RoPE (Su et al., 2024), which is used in LLMs like LLaMA (Touvron et al., 2023a, b) and Mistral (Jiang et al., 2023). Main strategies include embedding scaling (Chen et al., 2023; Liu et al., 2023b; Peng et al., 2023) and randomizing token positions (Ruoss et al., 2023; Zhu et al., 2023). Our emphasis is on the embedding scaling strategies.
Existing embedding scaling strategies adjust position embedding for longer sequences to match the pre-training range, avoiding feature extrapolation. For instance, Chen et al. (2023) compresses position indices to fit the pre-training range, extending LLaMA’s (Touvron et al., 2023a) context to 16K with 1,000 steps of fine-tuning. Alternatively, Liu et al. (2023b); Rozière et al. (2023); Xiong et al. (2023) modify RoPE’s rotary base and employ fine-tuning on extended sequences, termed Adjusted Base Frequency (ABF) or "NTK-aware" scaling. Code LLaMA (Rozière et al., 2023) achieved 16K context length with this method after 10,000 fine-tuning steps. YaRN (Peng et al., 2023) improved NTK-aware scaling by segmenting RoPE features and applying tailored extrapolation strategies, achieving 64K context length for LLaMA2 (Touvron et al., 2023b) with 400 fine-tuning steps. Distinguishingly, our Resonance RoPE focus on reducing feature interpolation on OOD positions, which we argue is another important factor in improving the length extrapolation capability of Transformer.
@@ 2.2 Long Context Evaluations
Evaluations of Transformer-based LLMs’ long-context capabilities are twofold: synthetic task assessments for length extrapolation strategies and real-world task evaluations at the LLM scale. Synthetic evaluations target simple tasks such as long sequence classification Tay et al. (2021) and arithmetic language modeling (Liu et al., 2023a; Kazemnejad et al., 2023). LLM scale evaluations measure metrics such as perplexity (PPL) in extensive text corpora (e.g., PG19 (Rae et al., 2020), GovReport (Huang et al., 2021), GitHub (Wu et al., 2022)) and complex tasks including summarization, question answering, and mathematical reasoning (An et al., 2023; Bai et al., 2023; Shaham et al., 2023).
@@ 3.1 Rotary Position Embedding (RoPE)
where ${\bm{W}}_{q},{\bm{W}}_{k}$ are trainable parameters, and $b$ is a constant called the rotary base, which is set to $10,000$  (Su et al., 2024) or other integers or fractions (Xiong et al., 2023; Peng et al., 2023). This form makes the dot product between the $m$-th query ${\bm{q}}_{m}$ and $n$-th key ${\bm{k}}_{n}$ only depend on the input ${\bm{x}}_{m},{\bm{x}}_{n}$ and their relative distance $(m-n)$:
@@ 3.2 Critical Dimensions of RoPE
In a TSTL scenario (Press et al., 2022), one takes a model trained on texts with lengths up to $L$, and tests it on a task with input lengths up to $L^{\prime}=sL$, with the scaling factor $s>1$. Recently, Liu et al. (2023b) discovered that there may exist two “critical dimensions” in RoPE features, which correspond to the dimensions $[2c:2c+1]$ that satisfies $\lambda_{c}\geq L$ and $\lambda_{c-1}<L$. The dimensions of RoPE features above and below the critical dimension (which we denote as “post-critical dimensions” and “pre-critical dimensions”, respectively) have different behaviors in TSTL: for post-critical dimensions (i.e., $j>c$), since their wavelengths satisfy $\lambda_{j}>L$, the training corpus does not cover all possible rotary angles $m\theta_{j}$ on a unit circle. Thus, these dimensions will encounter OOD value range on longer sequences. This is not an issue for pre-critical dimensions due to their shorter temporal wavelengths.
The concept of RoPE’s critical dimensions implicitly guides the development of RoPE scaling methods. For example, previous RoPE scaling methods (Chen et al., 2023; Xiong et al., 2023; Peng et al., 2023) mainly focus on reducing or avoiding value extrapolation on post-critical dimensions, and minimize post-training modifications to the pre-critical dimensions.
@@ 3.3 Yet another RoPE extensioN (YaRN)
YaRN (Peng et al., 2023) is the current state-of-the-art RoPE scaling method for TSTL. It introduces the “NTK-by-parts” scaling for RoPE, which applies different scaling strategies to each RoPE feature according to its temporal wavelength.
@@ 4 Proposed Method: Resonance RoPE
where $i=0,\dots,d-1$ and $\mathbb{X}\subset\mathbb{R}^{d}$ is the set of feature vectors to which we apply a position embedding.
Existing RoPE scaling methods (Xiong et al., 2023; Peng et al., 2023) mainly focus on the post-critical dimensions  of RoPE, since the rotary angle $m\theta_{j}$ on these dimensions extrapolates on OOD positions, hence creating a feature gap. In this section, we argue that reducing RoPE’s feature interpolation on the pre-critical dimensions  is also beneficial for better length extrapolation.
@@
Because of its simplicity, Resonance RoPE can be applied on top of RoPE and all RoPE-based scaling methods to reduce their feature gap in TSTL and further improve their performance. Meanwhile, this method only involves an offline computation of the scaled $\theta$, thus does not introduce online computation overhead.
@@ 5 Evaluating Position Embeddings with PosGen
Figure 2: An example of the three subtasks of PosGen. In this example, $h$ is a modular addition task with the modulus $m=7$ and the difficulty-controlling parameters $j=1,k=3$. The output token depends on: (1) only the local $j+k$ tokens in the recursive task; (2) $k$ local tokens and the beginning $j$ tokens in the CoT task; and (3) $k$ local tokens and $j$ tokens with a varied dependency distance in the semi-recursive task.
@@
Currently, neither perplexity-based evaluations (Rae et al., 2020; Huang et al., 2021; Wu et al., 2022) nor synthetic TSTL evaluations (Kazemnejad et al., 2023; Liu et al., 2023a) can effectively distinguish these two failure patterns, since the token generation difficulty tends to increase with respect to the sequence length in these tasks. To facilitate research on better position representations, we design PosGen, which controls the difficulty in generating tokens throughout the sequence to be identical, which effectively distinguishes the two types of TSTL failures. Failures in this benchmark are only due to the inability to recognize new token positions in TSTL scenarios.
@@
    Semi-recursive. This task simulates the token dependency pattern of the last-letter concatenation task (Zhou et al., 2023), where new tokens depend on both $k$ neighboring tokens (simulating the current progress) and $j$ tokens with varied distances (simulating the word sequence): $x_{l}=h(x_{\lfloor l-(j+k)/2\rfloor-j},\cdots,x_{\lfloor l-(j+k)/2\rfloor-1},\\ x_{l-k},\cdots,x_{l-1})$ when $l\geq j+k$.
Based on the equation for each subtask, when given the first $j+k$ tokens, one can generate a sequence with unlimited length as the ground truth sequence. We show an example of PosGen in Figure 2. As a TSTL benchmark, we train a model on a subtask with sequence length up to $L$, and evaluate the model’s accuracy on a longer sequence with length $L^{\prime}>L$ generated by the same rule on the unseen positions $L<m\leq L^{\prime}$, which we refer to as the “OOD Accuracy” (OOD Acc). This metric measures how well a model can recognize the OOD positions and continue following the generation rule learned during training. As a benchmark for position embeddings, a standard usage of this benchmark is to train a small Transformer (e.g., a 2-layer Transformer as used in our experiments) with different position embeddings on its training set with only short sequences, and test its OOD Accuracy on the test set with longer sequences. We provide our experiment setting for PosGen in more details in Section 6.1.1 and Appendix B.1.
@@ 6.1.1 Experiment Setup
We first apply Resonance RoPE on RoPE and YaRN, assessing the model’s performance on PosGen for unseen position recognition. We test on a modular addition task, which was proved to be learnable by a one-layer Transformer (Nanda et al., 2023). We configured $j=1,k=3$, and defined $h(x_{0},x_{1},x_{2},x_{3})=\sum_{i=0}^{3}{x_{i}}\mod 17$ with vocabulary $\mathbb{V}=\{0,\ldots,16\}$. Our experiments involved training a two-layer Transformer with different RoPE-based embeddings on sequences of length $L=64$, and evaluating on lengths of $L^{\prime}=256$ for OOD Accuracy. We generated 10,000 training sequences, and 1,000 each for validation and testing, and ensure that the first $j+k=4$ tokens in each sequence do not overlap to testify whether the model learns the correct generation mechanism. We averaged results over $5$ seeds. A more detailed setting is provided in Appendix B.1.
@@ 6.2.1 Experiment Setup
In this section, we apply our proposed Resonance RoPE to the current state-of-the-art RoPE scaling method, YaRN (Peng et al., 2023). More specifically, we replace the original position embeddings of LLaMA2 7B and 13B (Touvron et al., 2023b) with a series of scaled position embeddings, including the NTK-Aware scaling (bloc97, 2023; Xiong et al., 2023; Liu et al., 2023b), Dynamic NTK-Aware Scaling (Peng et al., 2023; Rozière et al., 2023), and YaRN (Peng et al., 2023).
For YaRN and Resonance YaRN, We use a scaling factor of $8$ and $4$ for LLaMA2 7B and 13B to extend their context window from $4$K to $32$K and $16$K, respectively. For the configurations that require fine-tuning, we fine-tune the LLM with the scaled position embedding on the training set of PG19 (Rae et al., 2020) with the fine-tuning setting and hyperparameters adopted directly from YaRN (Peng et al., 2023), with the only difference being that we control the total training token count to be approximately $100$M. A more detailed fine-tuning setting can be found in Appendix B.2. We test the model’s performance on two TSTL scenarios: language modeling evaluation on long-text sequences and long-text downstream application performance.
@@ 6.2.2 Perplexity on Long Sequence
We evaluate the model’s language modeling performance on GovReport (Huang et al., 2021) and Proofpile (Azerbayev, 2022). We randomly select $50$ samples from each dataset and report the final perplexity in text fragments of gradually increased length. We report the results in Figure 4. Of the tested methods, Resonance YaRN achieves the lowerst perplexity across all context lengths. Especially, Resonance YaRN achieves a lower perplexity compared to YaRN with the same set of hyperparameters optimized for YaRN, demonstrating the benefit of applying the Resonance technique to existing RoPE scaling methods.

### v2

@@ 1 Introduction
Canada CIFAR AI Chair. Corresponding author.
@@
Recent efforts to enhance TSTL performance have focused on LLMs equipped with Rotary Position Embedding (RoPE) (Su et al., 2024), such as LLaMA (Touvron et al., 2023a, b) and Mistral (Jiang et al., 2023), owing to their exceptional capabilities and widespread adoption. These initiatives aim to refine the test-time computation of RoPE position embedding by introducing a scaling factor to either the position index of each token (Chen et al., 2023) or RoPE’s base value (Xiong et al., 2023; Liu et al., 2024; Peng et al., 2024). These methods ensure that the position embeddings for out-of-distribution (OOD) positions remain within the range experienced during pre-training. This minimizes the need for the model to adapt to new position embedding value ranges, a task that is inherently difficult.
In this paper, we introduce Resonance RoPE, a novel technique designed to further narrow the generalization gap on position embeddings in TSTL scenarios. Recognizing that RoPE’s position embedding is governed by a complex, non-linear function, we posit that minimizing extrapolation on OOD positions, while crucial, is insufficient. We argue that it is equally vital to address the interpolation of RoPE features at the OOD positions. By implementing Resonance RoPE, we slightly scale each RoPE feature to correspond to an integer wavelength. This adjustment aligns each RoPE feature’s wavelength with a specific token span length, enabling it to "resonate" with a particular local context length. This simple modification effectively reduces the generalization gap for over half of the position embedding features in LLaMA and LLaMA2 under TSTL scenarios. Furthermore, our approach is compatible with RoPE and any RoPE-based scaling techniques, enhancing their performance in TSTL situations without the need for additional computational resources during training or inference.
Additionally, to facilitate further research on position embeddings, we present a new synthetic benchmark tailored for TSTL scenarios, named PosGen. Improving position embeddings for TSTL requires a detailed analysis of the cause of failures in handling longer contexts. However, current benchmarks, such as those measuring perplexity in long context (Rae et al., 2020; Huang et al., 2021; Wu et al., 2022) and most synthetic TSTL tasks (Liu et al., 2023; Kazemnejad et al., 2023) face a common issue: the difficulty of generating the next token increases with context length. This makes it difficult to determine whether a model’s failure is due to its inability to generate more complex tokens or its failure to recognize out-of-distribution (OOD) positions. PosGen addresses this limitation by standardizing the difficulty level of token generation across all positions. This ensures that any observed shortcomings are directly related to the model’s inability to identify and handle new token positions effectively.
@@ 2.1 Scaling of RoPE Position Encoding
Recent efforts in extending LLMs’ context window focus on manipulating position embedding (PE), particularly RoPE (Su et al., 2024), which is used in LLMs like LLaMA (Touvron et al., 2023a, b) and Mistral (Jiang et al., 2023). Main strategies include embedding scaling (Chen et al., 2023; Liu et al., 2024; Peng et al., 2024) and randomizing token positions (Ruoss et al., 2023; Zhu et al., 2024). Our emphasis is on the embedding scaling strategies.
Existing embedding scaling strategies adjust position embedding for longer sequences to match the pre-training range, avoiding feature extrapolation. For instance, Chen et al. (2023) compresses position indices to fit the pre-training range, extending LLaMA’s (Touvron et al., 2023a) context to 16K with 1,000 steps of fine-tuning. Alternatively, Liu et al. (2024); Rozière et al. (2023); Xiong et al. (2023) modify RoPE’s rotary base and employ fine-tuning on extended sequences, termed Adjusted Base Frequency (ABF) or "NTK-aware" scaling. Code LLaMA (Rozière et al., 2023) achieved 16K context length with this method after 10,000 fine-tuning steps. YaRN (Peng et al., 2024) improved NTK-aware scaling by segmenting RoPE features and applying tailored extrapolation strategies, achieving 64K context length for LLaMA2 (Touvron et al., 2023b) with 400 fine-tuning steps. Distinguishingly, our Resonance RoPE focus on reducing feature interpolation on OOD positions, which we argue is another important factor in improving the length extrapolation capability of Transformer.
@@ 2.2 Long Context Evaluations
Evaluations of Transformer-based LLMs’ long-context capabilities are twofold: synthetic task assessments for length extrapolation strategies and real-world task evaluations at the LLM scale. Synthetic evaluations target simple tasks such as long sequence classification Tay et al. (2021) and arithmetic language modeling (Liu et al., 2023; Kazemnejad et al., 2023). LLM scale evaluations measure metrics such as perplexity (PPL) in extensive text corpora (e.g., PG19 (Rae et al., 2020), GovReport (Huang et al., 2021), GitHub (Wu et al., 2022)) and complex tasks including summarization, question answering, and mathematical reasoning (An et al., 2023; Bai et al., 2023; Shaham et al., 2023).
@@ 3.1 Rotary Position Embedding (RoPE)
where ${\bm{W}}_{q},{\bm{W}}_{k}$ are trainable parameters, and $b$ is a constant called the rotary base, which is set to $10,000$  (Su et al., 2024) or other integers or fractions (Xiong et al., 2023; Peng et al., 2024). This form makes the dot product between the $m$-th query ${\bm{q}}_{m}$ and $n$-th key ${\bm{k}}_{n}$ only depend on the input ${\bm{x}}_{m},{\bm{x}}_{n}$ and their relative distance $(m-n)$:
@@ 3.2 Critical Dimensions of RoPE
In a TSTL scenario (Press et al., 2022), one takes a model trained on texts with lengths up to $L$, and tests it on a task with input lengths up to $L^{\prime}=sL$, with the scaling factor $s>1$. Recently, Liu et al. (2024) discovered that there may exist two “critical dimensions” in RoPE features, which correspond to the dimensions $[2c:2c+1]$ that satisfies $\lambda_{c}\geq L$ and $\lambda_{c-1}<L$. The dimensions of RoPE features above and below the critical dimension (which we denote as “post-critical dimensions” and “pre-critical dimensions”, respectively) have different behaviors in TSTL: for post-critical dimensions (i.e., $j>c$), since their wavelengths satisfy $\lambda_{j}>L$, the training corpus does not cover all possible rotary angles $m\theta_{j}$ on a unit circle. Thus, these dimensions will encounter OOD value range on longer sequences. This is not an issue for pre-critical dimensions due to their shorter temporal wavelengths.
The concept of RoPE’s critical dimensions implicitly guides the development of RoPE scaling methods. For example, previous RoPE scaling methods (Chen et al., 2023; Xiong et al., 2023; Peng et al., 2024) mainly focus on reducing or avoiding value extrapolation on post-critical dimensions, and minimize post-training modifications to the pre-critical dimensions.
@@ 3.3 Yet another RoPE extensioN (YaRN)
YaRN (Peng et al., 2024) is the current state-of-the-art RoPE scaling method for TSTL. It introduces the “NTK-by-parts” scaling for RoPE, which applies different scaling strategies to each RoPE feature according to its temporal wavelength.
@@ 4 Proposed Method: Resonance RoPE
where $i=0,\dots,d-1$ and $\mathbb{X}\subset\mathbb{R}^{d}$ is the set of feature vectors to which we apply a position embedding. Note that the formulation of the feature gap is similar to the “embedded vector distance” metric proposed by Xiong et al. (2023). However, these two metrics target totally different aspects of RoPE scaling methods. A more detailed comparison can be found in Appendix B.
Existing RoPE scaling methods (Xiong et al., 2023; Peng et al., 2024) mainly focus on the post-critical dimensions  of RoPE, since the rotary angle $m\theta_{j}$ on these dimensions extrapolates on OOD positions, hence creating a feature gap. In this section, we argue that reducing RoPE’s feature interpolation on the pre-critical dimensions  is also beneficial for better length extrapolation.
@@
Because of its simplicity, Resonance RoPE can be applied on top of RoPE and all RoPE-based scaling methods to reduce their feature gap in TSTL and further improve their performance. Meanwhile, this method only involves an offline computation of the scaled $\theta$, thus introducing no online computation overhead.
@@ 5 Evaluating Position Embeddings with PosGen
Figure 2: An example of the three subtasks of PosGen. This figure shows the process of generating the $12$th token shown in the red boxes for each subtask. In this example, $h$ is a modular addition task with the modulus $m=7$ and the difficulty-controlling parameters $j=1,k=3$. The output token depends on: (1) only the local $j+k$ tokens in the recursive task; (2) $k$ local tokens and the beginning $j$ tokens in the CoT task; and (3) $k$ local tokens and $j$ tokens with a varied dependency distance in the semi-recursive task.
@@
Currently, neither perplexity-based evaluations (Rae et al., 2020; Huang et al., 2021; Wu et al., 2022) nor synthetic TSTL evaluations (Kazemnejad et al., 2023; Liu et al., 2023) can effectively distinguish these two failure patterns, since the token generation difficulty tends to increase with respect to the sequence length in these tasks. To facilitate research on better position representations, we design PosGen, which controls the difficulty in generating tokens throughout the sequence to be identical, which effectively distinguishes the two types of TSTL failures. Failures in this benchmark are only due to the inability to recognize new token positions in TSTL scenarios.
@@
    Semi-recursive. This task simulates the token dependency pattern of the last-letter concatenation task (Zhou et al., 2023), where new tokens depend on both $k$ neighboring tokens (simulating the current progress) and $j$ tokens with varied distances according to a specific rule (simulating the word sequence): $x_{l}=h(x_{\lfloor l-(j+k)/2\rfloor-j},\cdots,x_{\lfloor l-(j+k)/2\rfloor-1},\\ x_{l-k},\cdots,x_{l-1})$ when $l\geq j+k$.
Based on the equation for each subtask, when given the first $j+k$ tokens, one can generate a sequence with unlimited length as the ground truth sequence. We show an example of PosGen in Figure 2. As a TSTL benchmark, we train a model on a subtask with sequence length up to $L$, and evaluate the model’s accuracy on a longer sequence with length $L^{\prime}>L$ generated by the same rule on the unseen positions $L<m\leq L^{\prime}$, which we refer to as the “OOD Accuracy” (OOD Acc). This metric measures how well a model can recognize the OOD positions and continue following the generation rule learned during training. As a benchmark for position embeddings, a standard usage of this benchmark is to train a small Transformer (e.g., a 2-layer Transformer as used in our experiments) with different position embeddings on its training set with only short sequences, and test its OOD Accuracy on the test set with longer sequences. We provide our experiment setting for PosGen in more details in Section 6.1.1 and Appendix C.1.
@@ 6.1.1 Experiment Setup
We first apply Resonance RoPE on RoPE and YaRN, assessing the model’s performance on PosGen for unseen position recognition. We test on a modular addition task, which was proved to be learnable by a one-layer Transformer (Nanda et al., 2023). We configured $j=1,k=3$, and defined $h(x_{0},x_{1},x_{2},x_{3})=\sum_{i=0}^{3}{x_{i}}\mod 17$ with vocabulary $\mathbb{V}=\{0,\ldots,16\}$.
Our experiments involved training a two-layer Transformer. Each layer follows T5-Small’s configurations (Raffel et al., 2020) except for the position embeddings. In this model, each attention head has $64$ dimensions. We apply different RoPE-based embeddings with the rotary base equal to $10,000$. The models are trained on sequences of length $L=64$, and evaluating on lengths of $L^{\prime}=256$ for OOD Accuracy. In this experiment setting, each head has $32$ RoPE features, out of which the first $17$ are pre-critical dimensions with a wavelength less than the maximum training length. We generated 10,000 training sequences, and 1,000 each for validation and testing, and ensured that the first $j+k=4$ tokens in each sequence do not overlap to testify whether the model learns the correct generation mechanism. We averaged results over $5$ seeds. A more detailed setting is provided in Appendix C.1.
@@ 6.2.1 Experiment Setup
In this section, we apply our proposed Resonance RoPE to the current state-of-the-art RoPE scaling method, YaRN (Peng et al., 2024). More specifically, we replace the original position embeddings of LLaMA2 7B and 13B (Touvron et al., 2023b) with a series of scaled position embeddings, including the NTK-Aware scaling (bloc97, 2023; Xiong et al., 2023; Liu et al., 2024), Dynamic NTK-Aware Scaling (Peng et al., 2024; Rozière et al., 2023), and YaRN (Peng et al., 2024).
For YaRN and Resonance YaRN, We use a scaling factor of $8$ and $4$ for LLaMA2 7B and 13B to extend their context window from $4$K to $32$K and $16$K, respectively. For the configurations that require fine-tuning, we fine-tune the LLM with the scaled position embedding on the training set of PG19 (Rae et al., 2020) with the fine-tuning setting and hyperparameters adopted directly from YaRN (Peng et al., 2024), with the only difference being that we control the total training token count to be approximately $100$M. A more detailed fine-tuning setting can be found in Appendix C.2. We test the model’s performance on two TSTL scenarios: language modeling evaluation on long-text sequences and long-text downstream application performance.
@@ 6.2.2 Perplexity on Long Sequence
We evaluate the model’s language modeling performance on GovReport (Huang et al., 2021) and Proofpile (Azerbayev, 2022). We randomly select $50$ samples from each dataset and report the final perplexity in text fragments of gradually increased length. We report the results in Figure 4. Of the tested methods, Resonance YaRN achieves the lowest perplexity across all context lengths. Especially, Resonance YaRN achieves a lower perplexity compared to YaRN with the same set of hyperparameters optimized for YaRN, demonstrating the benefit of applying the Resonance technique to existing RoPE scaling methods.